What is the wavelenghth of elextromagnetic radiation having a frequency of 5.00x10^12 Hz?
1
Expert's answer
2012-12-06T09:14:58-0500
λf = c
λ = Wavelength in meters
f = frequency in Hz
c = the velocity of the wave motion = 3*10^8 m/s
So
λ=3*10^8/5*10^12=6*10^-5